OpenStudy (anonymous):

Solve this equation. 1/4q=-20

OpenStudy (anonymous):

1/4q=-20 multiply both sides by 4 q = -20*4 q = -?

OpenStudy (texaschic101):

or....1/4q = - 20 (divide both sides by 1/4) 1/4q/1/4 = - 20/(1/4) q = -20 * 4/1 = (can you figure the answer from here ? )